# Stratified Selective Sampling for Instruction Tuning with Dedicated Scoring Strategy

### v1

@@ A.1 Instruction Categorization – details
Table 3: Training data overview for SetFit classifier.
@@ A.2 Difficulty Scorer – details
A.2.1 Correlations of difficulty/complexity metrics with input length
Figure 7: Relation between complexity/difficulty scores and length of the input sequence. The length of the input sequence should, in most cases, be unrelated to the difficulty of resolving it. Any correlation is therefore spurious.
As the length of an instruction might in some cases reflect the difficulty of the datapoint, there is no causal relationship between the two. A scorer that is trained to predict difficulty should therefore not rely on input length as a feature to rely its prediction upon. We show in Figure 7 how deita complexity is strongly correlated with input length, while the difficulty scorer we present in this paper is not. While this is not conclusive evidence that the complexity scorer is relying on input length as a feature in its prediction, these results are an indicator that it might. Further investigation is required to determine whether this relationship is causal.
A.2.2 Data generation difficulty scorer – details
@@
Figure 8 show the proportions of different instruction-response pairs that we use a training data for our difficulty scorer. The subsequent Table 4 shows the respective data source from which we obtained the data, as well as the type of evaluation metric that we use to evaluate the 18 LLMs from Table 5.
Figure 8: Setfit proportions of training data for difficulty scorer
@@
OpenBookQA (Mihaylov et al., 2018) Default train 299 acc
@@
Model Type Size
SmolLM2-135M-Instruct (Allal et al., 2025) Univ. 135M
SmolLM2-360M-Instruct (Allal et al., 2025) Univ. 360M
Qwen2.5-0.5B-Instruct (Team, 2024a) Univ. 0.5B
Qwen2.5-Math-1.5B-Instruct (Yang et al., 2024) Math 1.5B
Qwen2.5-Coder-1.5B-Instruct (Hui et al., 2024) Code 1.5B
Qwen2.5-1.5B-Instruct (Team, 2024a) Univ. 1.5B
SmolLM2-1.7B-Instruct (Allal et al., 2025) Univ. 1.7B
Qwen2.5-Math-7B-Instruct (Yang et al., 2024) Math 7B
Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) Code 7B
Qwen2.5-7B-Instruct (Team, 2024a) Univ. 7B
Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) Univ. 7B
Qwen2.5-14B-Instruct (Team, 2024a) Univ. 14B
Mistral-Small-24B-Instruct-2501 (Mistral AI, 2025) Univ. 24B
Qwen2.5-32B-Instruct (Team, 2024a) Univ. 32B
Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) Code 32B
Qwen3-32B (Team, 2025) Univ. 32B
gpt-4o-mini-2024-07-18 (Hurst et al., 2024) Univ. unk
gpt-4o-2024-08-06 (Hurst et al., 2024) Univ. unk
During the data collection for the difficulty scorer, we collect training sets from different benchmarks as well as data points from OpenAssistent (Köpf et al., 2023) for examples of open-ended generation. We evaluate these samples using the existing evaluation frameworks LM-evaluation harness (Gao et al., 2024), big-code evaluation harness (Ben Allal et al., 2022) and FastChat (Zheng et al., 2023b). For the LLM-as-a-judge evaluation, we use GPT4o as a judge and evaluate only a subset of 4 LLMs from Table 5 that we deem representative in terms of capabilities (gpt-4o-2024-08-06, Mistral-Small-24B-Instruct-2501, Mistral-7B-Instruct-v0.3 and SmolLM2-1.7B-Instruct).
A.2.3 Training the difficulty scorer – details
We equip regular CausalLLMs with a regression head by pooling the final dense layers and adding a linear projection to a scalar output, and finetune these models on the difficulty scoring problem. During finetuning, we use the hyperparameters detailed in Table 6. We finetune four different base models as difficulty scorer (Llama-3.1-8B (Grattafiori et al., 2024a), Qwen-2.5-7B (Team, 2024a), Qwen3-4B and Qwen3-8B (Team, 2025)) and converge on Qwen3-8B as it produces the best performance on our i.i.d. evaluation data.
Hyperparameter Value
batch_size 1
gradient_accumulation 16
learning_rate 1e-5
lr_scheduler_type linear
num_train_epochs 8
warmup_steps 100
max_seq_length 2048
weight_decay 0.01
neftune_noise_alpha 10
Table 6: Hyperparameter details for training the difficulty scorer
A.3 Code-quality Scorer – details
A.4 Instruction-following Scorer – details
Prompts used by LLM-annotator/judge for deriving instruction-following scores can be found in Figure 14, 15 and 16. As our test dataset, we collected responses from 10 LLMs on the IFEval benchmark (see Table 8), available on open-llm-leaderboard’s dataset collection of evaluation details. [^8] https://huggingface.co/open-llm-leaderboard Table 9 presents the evaluation results of various LLMs as the annotator/judge.
Model IFEval score
meta-llama/Llama-3.3-70B-Instruct 0.90
Qwen/Qwen2.5-14B-Instruct-1M 0.84
allenai/Llama-3.1-Tulu-3-70B-SFT 0.81
tiiuae/Falcon3-7B-Instruct 0.76
ibm-granite/granite-3.1-8b-instruct 0.72
microsoft/Phi-3-medium-128k-instruct 0.60
abacusai/Smaug-34B-v0.1 0.50
Qwen/Qwen2.5-32B 0.41
google/gemma-1.1-2b-it 0.31
databricks/dolly-v2-7b 0.20
Table 8: Performance of 10 considered LLMs on the IFEval benchmark.
Figure 9: Category proportions of IT datasets used in the experiments.
A.5 Datasets – details
Figure 9 shows SetFit classification results for IT datasets used in the experiments. Whereas Figure 10 shows the composition of source datasets for various sampling strategies. Deita is basically dominated by only two datasets: microsoft/orca-agentinstruct-1M-v1 (200k sampled) and WizardLMTeam/WizardLM_evol_instruct_V2_196K.
A.6 Finetuning – details
Hyperparameter falcon-10B llama-8B mistral-7B qwen-3B smollm-1.7B batch_size 4 8 8 8 8 gradient_accumulation 16 8 8 8 8 learning_rate 2.0e-05 1.0e-05 5.0e-06 5.0e-06 5.0e-06 lr_scheduler_type cosine num_train_epochs 2 attn_implementation flash_attention_2 warmup_ratio 0.03 max_seq_length 2048 weight_decay 0.1 neftune_noise_alpha 5 use_liger True
Table 10: Hyperparameter details for finetuning.
A.7 Evaluation – details
In the following, we list the benchmarks and the respective evaluation details, such as few-shot settings and other information about formatting and response extraction.
Benchmark Category ($\mathcal{L}$) Evaluation details IFeval (Zhou et al., 2023b) Generation 0-shot; prompt level strict acc GPQA (Rein et al., 2024) Multiple 0-shot BBH (Suzgun et al., 2022) Reasoning 3-shot; multiple-choice MuSR (Sprague et al., 2024) Reasoning 0-shot; multiple-choice Math (Hendrycks et al., 2021c) Math 4-shot ARC-C (Clark et al., 2018) Multiple 0-shot; multiple-choice GSM8k (Cobbe et al., 2021) Math 5-shot Hellaswag (Zellers et al., 2019) Multiple 0-shot; multiple-choice MMLU (Hendrycks et al., 2021b) Factual QA 0-shot; multiple-choice TruthfulQA (Lin et al., 2021) Factual QA 0-shot; multiple-choice Winogrande (Keisuke et al., 2019) Multiple 0-shot; multiple-choice MBPP (Austin et al., 2021) Coding 3-shot; pass@1 OpenBookQA (Mihaylov et al., 2018) Extraction 0-shot; multiple-choice
Table 11: List of used benchmarks with associated category and evaluation details.
A.8 Results – details
A.8.1 Results – scaling by benchmark
We present here the results of our scaling experiments, without aggregating them across evaluation benchmarks.
(a) ARC-challenge (Clark et al., 2018)
(b) GSM8k (Cobbe et al., 2021)
(c) HellaSwag (Zellers et al., 2019)
(d) BBH (Suzgun et al., 2022)
(e) GPQA (Rein et al., 2024)
(f) IFeval (Zhou et al., 2023b)
(g) Math (Hendrycks et al., 2021c)
(h) MuSR (Sprague et al., 2024)
(i) MBPP (Austin et al., 2021)
(j) MMLU (Hendrycks et al., 2021b)
(k) TruthfulQA (Lin et al., 2021)
(l) Winogrande (Keisuke et al., 2019)
Figure 11: Results scaling across all benchmark datasets
@@ A.8.1 Results – scaling by benchmark
5. Provide your answer in JSON format, with "review", "final_verdict", "code_original" and "code_revision" as keys.
@@ 2. Provide the expressed constraints in JSON format with the expressed constraint as the key
Answer the following questions. Provide the answer in JSON format with the question number as the key and the answer as
the value (true/false).
@@
Given the USER’s prompt and ASSISTANT’s response below, analyze whether the response addresses USER’s intents properly,
while respecting any constraints expressed in the prompt. Based on these judgments, provide the final score of response
quality in the range of 1 to 10, in JSON format (’score’ as key and quality score as value).

### v2

@@ A.2 Difficulty Scorer – details
A.2.1 Generating $\mathcal{D}_{\text{diff}}$
@@
Figure 8 shows the proportions of different instruction-response pairs that we use as the training data for the difficulty scorer. In Table 13 we report the benchmark data sources from which we obtained the data, as well as the type of evaluation metric that we use to evaluate the 18 LLMs from Table 4.
Figure 8: Proportions of different categories in the difficulty scorer training set $\mathcal{D}_{\text{diff}}$ after cleaning.
Model Type Size
SmolLM2-135M-Instruct (Allal et al., 2025) Univ. 135M
SmolLM2-360M-Instruct (Allal et al., 2025) Univ. 360M
Qwen2.5-0.5B-Instruct (Team, 2024a) Univ. 0.5B
Qwen2.5-Math-1.5B-Instruct (Yang et al., 2024) Math 1.5B
Qwen2.5-Coder-1.5B-Instruct (Hui et al., 2024) Code 1.5B
Qwen2.5-1.5B-Instruct (Team, 2024a) Univ. 1.5B
SmolLM2-1.7B-Instruct (Allal et al., 2025) Univ. 1.7B
Qwen2.5-Math-7B-Instruct (Yang et al., 2024) Math 7B
Qwen2.5-Coder-7B-Instruct (Hui et al., 2024) Code 7B
Qwen2.5-7B-Instruct (Team, 2024a) Univ. 7B
Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) Univ. 7B
Qwen2.5-14B-Instruct (Team, 2024a) Univ. 14B
Mistral-Small-24B-Instruct-2501 (Mistral AI, 2025) Univ. 24B
Qwen2.5-32B-Instruct (Team, 2024a) Univ. 32B
Qwen2.5-Coder-32B-Instruct (Hui et al., 2024) Code 32B
Qwen3-32B (Team, 2025) Univ. 32B
gpt-4o-mini-2024-07-18 (Hurst et al., 2024) Univ. unk
gpt-4o-2024-08-06 (Hurst et al., 2024) Univ. unk
During the data collection for the difficulty scorer, we collect training sets from different benchmarks as well as data points from OpenAssistent (Köpf et al., 2023) as samples of open-ended generation. We evaluate these samples using the existing evaluation frameworks LM-evaluation harness (Gao et al., 2024), big-code evaluation harness (Ben Allal et al., 2022) and FastChat (Zheng et al., 2023b). For the LLM-as-a-judge evaluation, we use GPT4o as the judge and evaluate only a subset of 4 LLMs from Table 4 that we deem representative in terms of capabilities (gpt-4o-2024-08-06, Mistral-Small-24B-Instruct-2501, Mistral-7B-Instruct-v0.3 and SmolLM2-1.7B-Instruct).
Figure 9 demonstrates the preprocessing step of calculating the deviation from the average as mentioned in Section 3.2.1.
Figure 9: An illustration of calculating the deviation of the average per dataset: Assuming we have a strong and a weak model with the average scores of $\bar{s}_{\text{s}}=0.8$ and $\bar{s}_{\text{w}}=0.15$ respectively. We subtract these averages from the respective item scores (here two examples with $s^{+}=1$ and $s^{-}=0$). We can see how an incorrect response of the strong model produces a much more negative score than an incorrect response from the weak model (compare $s_{\text{s}}^{-}$ and $s_{\text{w}}^{-}$). In other words, when a strong model gets an item wrong, this is much more meaningful for its difficulty compared to the weak model. Similarly, when a weak model correctly solves a datapoint, it is a much more meaningful signal for the data point’s ease than when a strong model is correct.
A.2.2 Difficulty scorer training
We equip regular CausalLLMs with a regression head by pooling the final hidden states and adding a linear projection to a scalar output, then finetune these models on the difficulty scoring task using the hyperparameters detailed in Table 5. We evaluate four different base models as difficulty scorer—Llama-3.1-8B (Grattafiori et al., 2024), Qwen-2.5-7B (Team, 2024a), Qwen3-4B and Qwen3-8B (Team, 2025))—and select Qwen3-8B, which achieves the best performance on in-distribution evaluation data.
Hyperparameter Value
batch_size 1
gradient_accumulation 16
learning_rate 1e-5
lr_scheduler_type linear
num_train_epochs 8
warmup_steps 100
max_seq_length 2048
weight_decay 0.01
neftune_noise_alpha 10
Table 5: Hyperparameter details for training the difficulty scorer
A.2.3 Effectiveness
To evaluate the effectiveness of the difficulty scorer for data selection across benchmarks, we sample 25k data points from the datasets in Table 2, either at random or based on Laser difficulty score ranking. The results in Figure 10(a) show that difficulty-based sampling outperforms random sampling on all benchmarks except GPQA. For comparison, Figure 10(b) reports results using the Deita complexity scorer under the same setup.
Figure 10: (a) Performance comparison of difficulty scorer vs. random sampling; (b) Performance comparison of complexity scorer vs. random sampling
Figure 11: Relation between complexity/difficulty scores and length of the input sequence. The length of the input sequence should, in most cases, be unrelated to its difficulty.
Figure 12: Relation between complexity/difficulty scores and the human code-forces difficulty scores. Difficulty scores show higher correlation to human scores than complexity scores.
[…]
A.2.4 Validity
We further evaluate whether the difficulty scores are akin to the notion of difficulty in humans or whether our scorer fits a spurious feature of our training data.
We first test it against instruction length, a common spurious feature of difficulty/complexity scorers. While difficult instructions are plausibly often longer than easier instructions, there is no causal relationship between the length and difficulty. A scorer that is trained to predict difficulty should therefore not rely on input length as a feature to rely its prediction upon. We show in Figure 11 how our difficulty scorer is weakly related to the spurious feature of instruction length, while the correlation of Deita complexity is much higher.
Besides showing how the difficulty scorer does not rely on instruction length, we also use an external criterion to evaluate whether it measures difficulty in the human sense. For this end, we score the CodeForces section of DeepMind’s code-contests dataset (Li et al., 2022) with our difficulty scorer as well as the Deita complexity scorer and correlate the scores with the given cf_rating. Figure 12 shows how the difficulty scorer is much better at predicting human difficulty scores than the Deita complexity scorer.
A.2.5 Experimentation with IRT-based targets
We further experimented with difficulty targets grounded in psychometric methods called item response theory (IRT; see Lord, 1968; Brzezińska, 2020; Van der Linden, 2018; Cai et al., 2016). IRT provides a framework for estimating latent traits from observed responses and has recently been applied to the evaluation of large language models (see e.g. Lalor et al., 2016; Rodriguez et al., 2021; Vania et al., 2021; Zhuang et al., 2023; Polo et al., 2024a). In our setting, an IRT model estimates item-level parameters from correctness scores across the LLMs from Table 4. We fit such a model separately on each dataset in $\mathcal{D}_{\text{diff}}$ and use the resulting item difficulty parameters $\beta$ as targets for training our difficulty scorer. We find that the IRT-based difficulty scorer yields improved results on the correlations with the Codeforces ratings (the validity criterion; see Figure 13). At the same time, this IRT-based difficulty scorer does not yield better results when used for data selection. We see the use of IRT-based difficulty scorers as an exciting future research direction, as the possibilities to gain more detailed insights into data point features are great (e.g. by using 2PL or 3PL instead of 1PL models or by extending from unidimensional to multidimensional models).
A.3 Quality Scorers – details
A.3.1 Instruction-following Scorer
Prompts used by LLM-annotator/judge for deriving instruction-following scores can be found in Figure 21, A.7 and 23. As our test dataset, we collected responses from 10 LLMs on the IFEval benchmark (see Table 6), available on open-llm-leaderboard’s dataset collection of evaluation details. [^8] https://huggingface.co/open-llm-leaderboard Table 7 presents the evaluation results of various LLMs as the annotator/judge.
Model IFEval score meta-llama/Llama-3.3-70B-Instruct 0.90 Qwen/Qwen2.5-14B-Instruct-1M 0.84 allenai/Llama-3.1-Tulu-3-70B-SFT 0.81 tiiuae/Falcon3-7B-Instruct 0.76 ibm-granite/granite-3.1-8b-instruct 0.72 microsoft/Phi-3-medium-128k-instruct 0.60 abacusai/Smaug-34B-v0.1 0.50 Qwen/Qwen2.5-32B 0.41 google/gemma-1.1-2b-it 0.31 databricks/dolly-v2-7b 0.20
Table 6: Performance of 10 considered LLMs on IFEval.
A.3.2 Code-quality Scorer
Figure 14: Performance of quality scoring-based sampling on various benchmarks.
A.3.3 Effectiveness
We test the efficacy of Laser task-specific quality scorers against the general-purpose Deita quality scorer, by comparing the performance of Mistral-7B-v0.3 finetuned on 25k data points, either randomly selected from data in Table 5, or based on ranking by considered quality scorers. For each considered benchmark, we selected only samples from the relevant category as defined in Table 3. Figure 14 demonstrates the clear gain of process reward model (PRM) scoring for Math benchmarks, particularly GSM8K. For Coding benchmarks, code-quality scoring outperforms Deita quality scoring on HumanEval, but achieves similar results as random. IF-quality scoring surpasses both Deita quality scoring and random on Generation/Brainstorming benchmarks. For other categories (Reasoning, Factual QA and Extraction), Deita quality scoring provides noticeable benefits only on TruthfulQA, with minimal impact on other benchmarks.
A.4 Sampling – details
A.4.1 Effectiveness
In Section 3.3, we introduce a sampling technique that aims to improve the diversity of the target dataset $\mathcal{D^{\prime}}$. The sampling technique consists of two components: firstly, we stratify the sampling by the categories that we assigned using the instruction classifier $\pi_{c}$ (according to the proportions shown in Figure 2 [laser]) and, secondly, we cluster the datapoints in embedding space and restrict the sampling to one datapoint per cluster. We sample here once 25k data points with the first technique (which we will call random+) and once 25k data points with both methods (correspondingly, random++) from $\mathcal{D}$. The random in the method name indicates that we select samples otherwise randomly (hence, not using any scoring during sampling). We then train a Mistral-7B-v0.3-base model on the resulting datasets and compare the results with entirely random sampling in Figure 16.
A.4.2 Efficiency
In Table 9, we compare the sampling runtime of Laser with Deita, which iteratively builds $\mathcal{D^{\prime}}$ by adding dissimilar samples based on embedding distance, as well as CaR, which runs clustering on the entire dataset $\mathcal{D}$. While Deita is highly efficient given small $m$ (e.g., 1k), its runtime grows rapidly with larger $m$. Meanwhile, CaR suffers from out-of-memory issues at large $m$. In contrast, Laser remains stable across different $m$ and is generally more efficient than CaR.
$m$ Deita CaR Laser
1k 92.3 s 5515.7 s 1505.1 s
10k 970.2 s 5750.9 s 1555.9 s
100k 41055.3 s OOM 2818.1 s
Table 9: Runtime comparison for different $m$.
Figure 16: Effect of different sampling strategies on training outcomes. While random selects $m$ = 25k arbitrary samples, random+ samples by stratifying the categories in $D^{\prime}$ according to the proportions shown in Figure 2 (laser). Ultimately, random++ also includes clustering (as described in Section 3.3, while randomly selecting a datapoint from each cluster)
A.5 Datasets – details
A.5.1 Datasets main experiments
Figure 18 shows SetFit classification results for IT datasets from which we composed $\mathcal{D}$. On the other hand, Figure 15 shows which of the source datasets are sampled when we use various sampling strategies. Interestingly, while Laser maintains relative diversity across different source datasets, Deita is completely dominated by only two datasets: microsoft/orca-agentinstruct-1M-v1 (200k sampled) and WizardLMTeam/WizardLM_evol_instruct_V2_196K.
A.5.2 Datasets robustness experiments
Dataset details $\mathcal{D}_{\text{strong}}$ and $\mathcal{D}_{\text{weak}}$
In Section 4.2.2, we introduce $\mathcal{D}_{\text{strong}}$ and $\mathcal{D}_{\text{weak}}$. While $\mathcal{D}_{\text{strong}}$ simply consists of the Tülu v3 dataset, $\mathcal{D}_{\text{weak}}$ is a aggregation of different datasets. The single components exhibited either in our own pilot experiments or in prior related work comparably weak performance to other datasets. This weaker performance also shows, when we compare the finetuning using full $\mathcal{D}_{\text{weak}}$ with the finetuning using full $\mathcal{D}_{\text{strong}}$ in Figure 6, with the latter significantly outperforming the former.
Dataset #samples (#turns) ConiferLM/Conifer (Sun et al., 2024) 13k databricks/databricks-dolly-15k (Conover et al., 2023) 15k akoksal/LongForm (Köksal et al., 2023) 23k alpaca (Taori et al., 2023) 52K vicgalle/alpaca-gpt4 52K ai2-adapt-dev/flan_v2_converted (Sanh et al., 2022) 90K nvidia/Daring-Anteater (Wang et al., 2024) 100k AI-MO/NuminaMath-CoT999random subset (Li et al., 2024a) 250k all 595K
Table 10: Composition of $\mathcal{D}_{\text{weak}}$
Figure 17: Category distribution for datasets sampled from the 5 bias source datasets $\mathcal{D}_{\text{skewed}}$ using different sampling strategies.
Results of experiments with $\mathcal{D}_{\text{skewed}}$
In Section 4.2.2, we experiment with skewed data distributions. In Figure 17, we show the resulting category distributions when sampling with different sampling approaches from the 5 skewed source distributions. As we can see, only Laser reduces the introduced bias in the skewed source data, while Deita is not effective in reducing the bias, but in many cases exacerbates it.
A.6 Finetuning – details
Hyperparameter falcon-10B llama-8B mistral-7B qwen-3B smollm-1.7B
batch_size 4 8 8 8 32
gradient_accumulation 16 8 16 8 16
learning_rate 2.0e-05 5.0e-06 5.0e-06 2.0e-05 1.0e-04
num_train_epochs 2 2 2 2 3
weight_decay 0.1 0.01 0.01 0.01 0.01
warmup_ratio 0.03 0.03 0.1 0.03 0.03
lr_scheduler_type cosine
attn_implementation flash_attention_2
max_seq_length 2048
neftune_noise_alpha 5
use_liger True
Table 11: Hyperparameter details for finetuning.
A.7 Results – details
Figure 24 shows the scaling experiment results across benchmarks, reported without normalizing score gain.
Figure 18: Category distribution in IT datasets used in the experiments.
Table 12: Training data overview for SetFit classifier.
@@ A.7 Results – details
OpenBookQA (OpenBookQA2018) Default train 299 acc
@@
5. Provide your answer in JSON format, with r̈eview", "final_verdict", "code_original" and "code_revision" as keys.
@@ 2. Provide the expressed constraints in JSON format with the expressed constraint as the key
Answer the following questions. Provide the answer in JSON format
with the question number as the key and the answer as the value
(true/false).
@@
Given the USER’s prompt and ASSISTANT’s response below,
analyze whether the response addresses USER’s intents properly,
while respecting any constraints expressed in the prompt.
Based on these judgments, provide the final score of response
quality in the range of 1 to 10, in JSON format (’score’ as key
and quality score as value).
@@
(a) AGIEval (Zhong et al., 2024)
(b) AlpacaEval (Dubois et al., 2024)
(c) ARC-Challenge (Clark et al., 2018)
(d) BBH (Suzgun et al., 2023)
(e) GPQA (Rein et al., 2024)
(f) HellaSwag (Zellers et al., 2019)
[…]
(i) Math (Hendrycks et al., 2021c)
(j) HumanEval (Chen et al., 2021)
(k) MBPP (Austin et al., 2021)
(l) IFeval (Zhou et al., 2023b)
(m) MMLU (Hendrycks et al., 2021b)
(n) MuSR (Sprague et al., 2024)
(o) OpenBookQA (Mihaylov et al., 2018)
(p) TruthfulQA (Lin et al., 2022)
(q) Winogrande (Sakaguchi et al., 2021)
Figure 24: Results scaling across all benchmark datasets